how long did he have to hold his breath during the div

12. No-limits apnea is an AIDA International freediving discipline in which the freediver descends, at around 5 m/s, using a heavy metal bar or sled grasped by the diver and fixed to a line. They then ascend at around 2 m/s via inflatable lifting bags or vests with inflatable compartments. No-limits apnea involves diving to great depths, while holding their breaths, and has claimed the lives of several divers. The current no-limits world record holder is Herbert Nitsch with a depth of 253.2 m, who described a bewildering blackout through narcosis feeling during his dive and subsequently received extensive decompression treatment. How long did he have to hold his breath during the dive?
Answer #1

t = t1 + t2

t1 and t2 = time of descend and ascend

t1 = d/v1 = 253.2/5 = 50.64 s

t2 = d/v2 = 253.2/2 = 126.6 s

t = 177.24 sec